### v1

# Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning

Abstract
Large language models (LLMs) are rapidly changing various domains. However, their capabilities in handling conversational breakdowns still require an in-depth exploration. This paper addresses the challenge of detecting and mitigating dialogue breakdowns within LLM-driven conversational systems. While powerful models from OpenAI and Anthropic excel in many dialogue tasks, they can still produce incoherent or contradictory responses, commonly referred to as breakdowns, which undermine user trust. To tackle this, we propose an approach that combines specialized fine-tuning with advanced prompting strategies, including few-shot learning, chain-of-thought reasoning, and analogical prompting. In particular, we fine-tune a small 8B model and demonstrate its robust classification and calibration capabilities in English and Japanese dialogue. We also validate its generalization on the BETOLD dataset, achieving a 7% accuracy improvement over its base model. Furthermore, we introduce a real-time deployment architecture that selectively escalates suspicious responses to more resource-intensive frontier models only when breakdowns are detected, significantly cutting operational expenses and energy consumption. Experimental results show our method surpasses prior state-of-the-art specialized classifiers while also narrowing performance gaps between smaller open-source models and large proprietary ones. Our approach offers a scalable solution for robust conversational AI in high-impact domains by combining efficiency, interpretability, and reliability.
{IEEEImpStatement}
Our research advances conversational AI by reducing incoherent or contradictory responses, known as dialogue breakdowns, which erode user confidence. Unlike previous approaches dependent on large, specialized models, we achieve state-of-the-art accuracy and strong calibration using a fine-tuned 8B-parameter language model. This improvement significantly cuts operational costs while maintaining reliability in detecting breakdowns in multilingual and task-oriented dialogues. We further reduce compute demands by calling on larger, more resource-intensive models only when absolutely necessary. This balance between scalability, accuracy, and sustainability facilitates the greater adoption of safe, trustworthy AI systems in critical areas such as customer support, education, and healthcare. By aligning system confidence with human judgment, our framework promotes greater transparency and user acceptance, paving the way for more robust and accessible deployment of next-generation conversational technologies.
{IEEEkeywords}
large language models, conversational artificial intelligence, human-computer interaction
1 Introduction
Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3]. The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], emphasizes both the unique capabilities and the immense potential of LLMs. These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally redefining human-AI interactions [6, 7, 8].
Table 1: Sample Dialogue
         1.   Assistant: It’s nice to go shopping alone.
         2.   User: I agree. That’s nice.
         3.   Assistant: Shopping takes time.
         4.   User: Window shopping is also fun.
   Determine if the following utterance leads to a dialogue breakdown:
         5.   Assistant: It’s fun to go shopping with somebody.
Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13]. Table 1 illustrates an example scenario where the assistant’s initial and subsequent utterances contradict each other, illustrating a common breakdown scenario.
The importance of addressing dialogue breakdowns grows alongside adopting LLMs in high-stakes environments, such as customer support and educational applications [6, 7, 8]. Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15]. Consequently, robust methodologies are clearly needed to detect and mitigate dialogue breakdowns to preserve conversational quality and user trust.
Previous research in dialogue breakdown detection has predominantly focused on specialized classifiers trained on labeled datasets like the Dialogue Breakdown Detection Challenge [12, 16]. However, the generalization of these models to diverse, real-world contexts remains limited. Recent literature also indicates that generalist LLMs, despite their flexibility, still lag behind human-level performance in nuanced conversational tasks, indicating a persistent research gap[17, 18].
Addressing these challenges, we introduce innovative approaches that take advantage of generalist LLMs’ reasoning capabilities through supervised fine-tuning and structured prompting. Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by the superior Llama-3.3 70B model. These distilled reasoning tracks guide and enhance interpretability during training, significantly improving the model’s decision-making process. We rigorously evaluate our model’s generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities. We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27]. These strategies encourage systematic reasoning and adaptability without extensive labeled datasets.
Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge. Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability. Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary. Given that high-capacity models significantly raise operational costs and energy consumption, our architecture optimizes resource utilization, achieving substantial cost reductions and promoting sustainability.
Our contributions are summarized as follows:
  • Conducted the first systematic comparative analysis between open-source and frontier closed-source models, setting new benchmarks in dialogue breakdown detection.
  • Extensively evaluated model calibration alongside accuracy, revealing critical insights into reliability and overconfidence across diverse conversational scenarios.
  • Proposed an efficient, real-time deployment architecture that significantly reduces operational costs and promotes sustainability by selectively invoking large-scale models.
2 Related Work
Ensuring robustness in conversational AI systems, particularly in dialogue breakdown detection and mitigation, is a vital area of research. Dialogue breakdowns refer to situations where conversational coherence or relevance is disrupted, hindering the conversation’s smooth progression and diminishing user satisfaction [12, 13]. Effective conversational agents must not only detect these breakdowns but also address them to maintain user engagement and trust [28, 29].
2.1 Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section 5.1 for more information) [12]. The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification. For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32]. BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33]. This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
Further advances in semi-supervised learning have also significantly impacted the domain. The S2T2 model, for example, introduced a dual-teacher training paradigm to leverage both labeled and unlabeled dialogue data [34]. S2T2 employs two teacher models: one trained on high-quality labeled data and another trained on masked dialogues variations, to guide a student model collaboratively. This hybrid training strategy has achieved new state-of-the-art results on the DBDC5 dataset, surpassing earlier methods. Specifically, S2T2 utilizes RoBERTa-large for the English track and XLM-R-large (a multilingual variant of RoBERTa) combined with a context-matching mechanism for the Japanese track [35, 36, 34].
Recent studies have also explored the potential of general-purpose large language models (LLMs) in dialogue breakdown detection. Finch et al. [17, 18] evaluated ChatGPT’s performance in identifying dialogue behaviour across nine distinct categories defined in the ABC-Eval dataset. Their findings reveal that while ChatGPT shows promise, even outperforming specialized models in identifying empathetic behaviour, it still falls short of human performance in other dialogue tasks. This suggests that current large models still face challenges in reliably identifying breakdown instances despite their sophisticated conversational abilities and extensive world knowledge. Nonetheless, using LLMs for dialogue evaluation is a promising direction since they bring broad world knowledge and understanding of conversational dynamics. To our knowledge, no other work has evaluated LLMs for dialogue breakdown detection and remediation.
2.2 Current State of Conversational Agents
Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns. For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12]. Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2]. Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the chance of breakdowns due to factual errors or contradictions [37]. The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
2.3 Techniques in Conversational AI
To further improve the reliability of LLM conversational agents, researchers are exploring a variety of techniques:
Analogical Reasoning. This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem. Instead of relying solely on provided examples, the model generates analogous examples as context. Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38]. It ”prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25]. In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has seen and use that to form a better response to a user query.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (”thinking aloud”) before giving a final answer. This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24]. For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response. This reduces mistakes and non-sequiturs. CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly ”enhances their ability to reason” about the conversation [24, 40, 39].
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all. In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query. The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating ”language models as few-shot learners.” [22]. This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model. LLMs can often do this for dialogue tasks. For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning. In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
3 Problem Definition
Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29]. Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11]. As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount. Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in compounded user frustration [14, 15].
We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A). The dialogue proceeds in pairs of utterances, where at turn $i$, the user produces an utterance $u_{i}$, and the agent replies with $s_{i}$. Formally, we may represent the dialogue as follows:
$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}$
We aim to detect, at each agent utterance $s_{i}$, whether the conversation has experienced a breakdown in coherence, relevance, or consistency.
3.1 Utterance-Level Breakdown Detection
Let $\mathcal{H}_{i}$ denote the contextual history available just before the agent produces its $i$-th response:
$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}$
We define a classification function $f$ that, given $\mathcal{H}_{i}$ and the agent’s latest utterance $s_{i}$:
$f:\bigl{(}\mathcal{H}_{i},\;s_{i}\bigr{)}\mapsto(\hat{b}_{i},\;\hat{c}_{i},\;\hat{j}_{i}),$
where:
  • $\hat{b}_{i}\in\{0,1\}$ is a binary classification indicating dialogue breakdown (1) or non-breakdown (0).
  • $\hat{c}_{i}\in[0,1]$ is a confidence score representing the model’s certainty about the predicted label.
  • $\hat{j}_{i}$ is a textual justification explaining the model’s reasoning process.
For a complete dialogue $D$ of length $n$ system turns, the output $\mathcal{O}$ of the detection system is thus:
$\mathcal{O}(D)=\{(b_{i},c_{i},j_{i})\}_{i=1}^{n}$
3.2 Consolidation of Three-Class Annotations
Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB). Given the subjectivity of human annotations in dialogue breakdown detection, we often consolidate PB into B. When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB. We can convert each utterance into a binary label $b_{i}$ by thresholding:
$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
here $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, “breakdown vs. non-breakdown” refers to this binary mapping.
3.3 Conversation-Level Labeling
In task-oriented datasets such as BETOLD [19], the focus is on the entire conversation outcome. Each dialogue $D$ is labeled a failure if, for instance, the user hangs up or requests escalation to a human agent. We denote this conversation-level label as
$\mathcal{O}(D)\;=\;\begin{cases}1,&\text{if the conversation leads to breakdown},\\ 0,&\text{otherwise}.\end{cases}$
In this case, the model seeks to predict after observing all $(u_{i},s_{i})$ pairs in $D$. This differs from utterance-level detection since any local breakdown event might cause the entire conversation to fail.
3.4 Confidence Calibration
An effective breakdown detector should classify accurately and calibrate its confidence well. Let $p_{i}$ be the true probability of a breakdown for the $i$-th utterance (estimated from multiple annotators) and $\hat{c}_{i}$ be the model’s predicted probability. We measure calibration quality via mean-squared error (MSE):
$\mathrm{MSE}\;=\;\frac{1}{N}\sum_{i=1}^{N}\Bigl{(}\,\hat{c}_{i}\;-\;p_{i}\Bigr{)}^{2},$
where $N$ is the total number of utterances in the test set. Lower MSE indicates that the model’s self-reported confidence aligns more closely with actual annotator distributions.
4 Methodology
Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations. We focus on three key components: a compact fine-tuned model to recognize dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and propose a multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
4.1 Supervised Fine-tuning with Reasoning Augmentation
To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT). We choose Llama-3.1 8B [3] as a balance between accuracy and computational cost. The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels. Let:
  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ be the training data, where each sample has context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
  • $T$ be a larger “teacher” LLM, such as Llama-3.3 70B that can generate synthetic reasoning traces $r_{i}$ (i.e., a chain-of-thought explanation) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$. We fine-tune our student model $S$ to minimize the cross-entropy loss ($\mathcal{L}_{CE}$) [41] between predicted and true labels:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})],$
here, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown. The phase allows the student to see both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
4.2 Advanced Prompting Strategies
While a fine-tuned compact model is efficient for real-time monitoring, powerful but costly LLMs can still be employed through strategic prompt engineering. Let $\mathcal{G}$ denote a generalist LLM, for instance GPT-4 or DeepSeek-R1. Given a conversation snippet $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under a chosen strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$. We then parse $g(\Pi(\mathcal{H}_{i},s_{i};\,\alpha))$ to obtain a breakdown label and confidence. The main prompting strategies we explore are:
Zero-Shot (ZS) Prompting. We provide only a task description and the current example $(\mathcal{H}_{i},s_{i})$. The model must infer the classification criterion from the instructions alone. A sample zero-shot prompt for DBDC5 is shown in Figure 1.
Figure 1: Sample Zero-Shot Prompt for DBDC5
Few-Shot (FS) Prompting. We supply $k$ labeled examples of breakdown vs. non-breakdown before the new sample. Formally,
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i})$
Within the FS category, we differentiate:
  • 2-Shot Easy (2S-Easy): Two clear dialogues (one breakdown and one non-breakdown) where human annotators were very confident in the labels (>80% agreement). For example, an easy non-breakdown case might be a smooth dialogue that successfully completes, and an easy breakdown case might feature an obvious user hang-up after a system error. This helps the model anchor on unambiguous prototypes.
  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%). One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low). We hypothesize exposing the model to ambiguous scenarios can improve its performance under uncertainty, hopefully teaching it to output more calibrated confidence in similar situations.
  • 4-Shot (4S): We combine the Easy and Hard exemplars—two from 2S-Easy and two from 2S-Hard—thus exposing the model to a broader range of scenarios. While 4S can yield higher accuracy, prompt length grows, which may overwhelm smaller models.
    In datasets without direct human annotations such as BETOLD, we select short (15-20 turns) vs. long (21-30) dialogues as “easier” vs. “harder” exemplars, then mimic the above FS setups accordingly. In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
4.3 Self-Guided Reasoning Techniques
To further enhance the model’s reasoning depth and generalizability, we used and proposed advanced prompting techniques designed to encourage structured, self-generated reasoning:
Chain-of-Thought (CoT). We encourage multi-step reasoning by instructing $g$ to “think step-by-step” before deciding:
[…]
This often yields more coherent predictions on complex queries, albeit at higher token usage.
Analogical Reasoning (AR). Instead of providing examples directly, we ask $\mathcal{G}$ to generate hypothetical analogous conversations $\mathcal{A}_{i}$ from its internal knowledge, then classify the current conversation. Formally:
$\mathcal{A}_{i}\;=\;\mathcal{G}\!\bigl{(}\Pi_{\mathrm{AR}}(\mathcal{H}_{i},s_{i})\bigr{)},$
where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$. The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$. This self-guided technique removes the need for hand-crafted exemplars, potentially offering more tailored guidance to each test case.
4.4 Our Proposed Reasoning Technique
Curricular Learning with Analogical Reasoning (CL+AR). Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn. The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks. We propose applying this idea to LLM reasoning: the model is first steered to solve easy ”proxy” queries related to the target query, then gradually given harder versions of the problem. The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought. In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity). This gradually “walks” the model from simpler to harder analogies. We then request the final decision for $(\mathcal{H}_{i},s_{i})$. Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies. The overall prompt is:
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}$
We find that this method improves classification on challenging dialogues, although it may increase output token length.
4.5 Deployment Architecture
While a powerful LLM can reliably detect breakdowns, frequent usage may be cost-prohibitive. We address this via a real-time, hierarchical system (Figure 2) with three main modules to ensure dialogue robustness and maintain user trust:
  1. 1.
    AI Assistant: Generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and context $\mathcal{H}_{i}$:
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
    Here, $\mathcal{G}_{\mathrm{assistant}}$ can be a moderately large language model such as Llama-3.3 70B.
  2. 2.
    Dialogue Disruption Monitor: Before presenting the assistant-generated response $s_{i}$ to the user, our fine-tuned model rapidly assesses the response for potential breakdowns or unsafe interactions, outputting:
    $(\hat{b}_{i},\;\hat{c}_{i},\;\hat{j}_{i})\;=\;\mathcal{G}_{\mathrm{monitor}}\!\bigl{(}\mathcal{H}_{i},\;s_{i}\bigr{)},$
    where $\hat{b}_{i}\in\{0,1\}$ is the predicted breakdown label, $\hat{c}_{i}\in[0,1]$ is the confidence, and $\hat{j}_{i}$ (optional) is a justification. If $\hat{b}_{i}=0$ or $\hat{c}_{i}<T$, $s_{i}$ is deemed acceptable. Otherwise, we escalate to a larger “superior” LLM. The sub-second detection latency supports seamless user experiences.
  3. 3.
    Superior Model: Upon intervention, this model revises the response to prevent breakdown and ensure user safety:
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
    The monitor then re-checks $\bigl{(}\mathcal{H}_{i},\;s_{i}^{*}\bigr{)}$. This process can iterate up to $k$ times or until the breakdown is resolved.
This multi-tiered strategy significantly reduces computational demands and ensures the most resource-intensive models are only utilized in critical scenarios, thereby enhancing both efficiency and user experience in conversational AI deployments.
Figure 2: Real-time Response Correction Architecture. The dialogue disruption monitor intercepts potentially unsafe assistant responses, triggering a correction from a superior model before presenting the response to the user.
5 Experiments
We evaluate our proposed approach using three prominent dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
5.1 Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12]. The test set contains annotated dialogues in both English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents. Each dialogue turn is annotated by 15–30 human annotators, classifying system utterances into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty). DBDC5 (and its predecessors DBDC3/4) have been important for research in identifying when dialogues go off-track and why.
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19]. This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a ”late user-initiated hang-up or forward” (LUHF). LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention. Approximately 33% of the dataset represents breakdown cases. BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy. We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
5.2 Data Preprocessing and Experimental Setup
Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown). Ambiguous annotations (’Possible Breakdown’) are consolidated into the Breakdown class. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
Our evaluation requires the LLM to provide three fields per response: justification, decision, and confidence score. Experimentally, the ordering of these requests influenced the model’s performance significantly. When the model was first prompted for a decision, subsequent justifications often appeared overly confident and less reflective. Conversely, requesting justifications first yielded more nuanced and deliberative decisions, enhancing overall output quality.
For advanced prompting methods (AR and CL+AR), we conducted experiments on a random subset representing 10% of each dataset due to higher token demands. Ablation studies indicated that a two-pass analogical reasoning approach (initial analogy generation followed by integration) performed comparably to single-pass prompting, aligning with findings in existing literature [25].
5.3 Fine-tuning Parameters
We fine-tune a Llama-3.1 8B instruct model using Low-Rank Adaptation (LoRA, rank=16) [42], employing an AdamW 8-bit optimizer [43] with a learning rate of $2\times 10^{-4}$, linear decay, batch size of 8, and weight decay of 0.01 across 3 epochs. Training utilized a single NVIDIA A100 40GB GPU. We hypothesize that a finely tuned smaller LLM can balance competitive accuracy with efficiency, making it viable for real-time dialogue monitoring. Our trained dialogue disruption monitor is openly available on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor, facilitating community engagement. Emphasizing sustainability, our fine-tuning approach, which leverages LoRA’s partial weight updates, significantly reduces computational load and energy consumption compared to traditional full-scale training.
5.4 Evaluation Metrics
We use accuracy and F1 score as our primary evaluation metrics for breakdown detection. Accuracy measures the overall correctness of breakdown vs non-breakdown predictions. F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class. We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34]. This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results. Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits). We parse the model’s JSON-formatted output to extract the decision label. In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to verify output alignment with ground-truth labels [44].
On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE of the annotators’ probability distribution for a dialogue and the LLM’s verbalized confidence in its decision.
5.5 Inference and Cost Considerations
For a comprehensive study, we examine a diverse range of LLMs, as listed in Table 2. The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4) [1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2) [2]. In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B) [20], DeepSeek (DeepSeek-R1) [21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B) [3] offer greater transparency to model details, increased flexibility for customization, and are often more cost-effective. For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints. This approach allowed us to evaluate different models under a consistent framework and logging. However, using these hosted models incurs costs per token. In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens. We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage. Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off. They may improve reasoning but use more tokens, impacting cost and latency. All models use a temperature of 0, which controls the randomness of language model output; a top-p value of 0.9, known as nucleus sampling, which further influences randomness; and a maximum token limit of 2048, reflecting the context-length constraints of smaller LLMs such as Mixtral 8x7B and certain Llama variants.
6 Results and Discussion
This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks). We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor. Table 2 summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
Family Model Prompt BETOLD DBDC5 English DBDC5 Japanese
[…]
Previous SOTA BERT+SSMBA 73.9 78.2
S2T2 77.9 82.4 76.7 75.4
Anthropic Claude-3.5 Haiku ZS 74.4 55.0 82.1 80.4 86.3 65.3 5.9 74.6 76.0 73.0 7.3
CoT 73.9 52.6 82.0 82.0 86.9 71.2 6.8 77.1 76.5 77.6 8.2
2S (Easy) 74.0 64.3 79.5 82.5 87.5 70.9 5.5 68.9 73.3 62.9 10.6
2S (Hard) 74.1 66.9 78.8 82.5 87.4 70.9 4.1 66.3 72.0 57.7 8.4
4S 74.1 67.0 78.7 82.9 87.6 72.5 4.4 67.0 72.3 59.2 10.2
AR 76.3 68.6 81.0 78.5 85.8 55.7 82.0 88.5 59.1
CL+AR 77.0 64.4 83.1 78.0 85.4 55.1 78.0 85.9 50.0
Claude-3.5 Sonnet v2 ZS 75.1 54.9 82.8 82.7 87.5 71.9 8.1 81.3 81.2 81.4 8.2
CoT 76.6 56.7 84.0 82.5 87.3 71.6 7.8 78.8 78.7 78.9 8.3
2S (Easy) 75.8 60.4 82.5 83.5 88.2 72.3 7.1 74.0 76.7 70.5 10.8
2S (Hard) 76.9 62.9 83.3 81.2 86.9 66.7 7.5 71.2 75.3 65.6 10.3
4S 76.7 64.4 82.7 84.0 88.7 72.7 6.5 71.4 75.4 65.9 10.7
AR 73.3 55.0 81.1 85.5 89.8 74.8 88.0 91.7 78.6
CL+AR 76.3 63.6 82.4 83.5 88.5 70.8 89.0 92.4 80.0
OpenAI GPT-3.5 Turbo ZS 41.2 51.4 25.8 68.7 80.5 21.2 16.8 50.8 64.4 20.4 25.2
CoT 43.1 51.8 30.6 67.9 77.9 41.9 17.8 55.1 58.9 50.5 22.0
2S (Easy) 64.5 62.3 66.5 67.2 76.6 44.9 14.7 56.5 61.8 49.6 21.5
2S (Hard) 57.3 52.9 60.9 67.0 74.7 52.8 8.6 51.1 64.6 21.0 16.0
4S 47.5 55.0 37.1 70.0 77.1 56.4 12.8 57.1 57.7 56.4 20.5
AR 43.0 52.2 29.4 71.5 81.9 32.9 71.0 82.2 21.6
CL+AR 42.2 51.9 27.8 68.0 79.2 30.4 72.0 81.8 39.1
GPT-4o ZS 74.1 47.8 82.7 81.4 85.9 72.5 9.2 79.2 76.7 81.2 9.8
CoT 73.2 43.6 82.5 82.3 86.5 74.3 9.1 79.5 77.6 81.2 9.8
2S (Easy) 75.6 53.3 83.5 82.7 86.7 75.2 7.2 79.3 78.9 79.6 8.7
2S (Hard) 77.4 63.0 83.7 83.5 87.3 76.3 5.1 80.2 77.7 82.1 6.4
4S 77.7 62.8 84.1 82.0 86.1 74.6 5.8 79.8 79.2 80.4 7.5
AR 70.4 42.9 80.0 80.5 86.3 66.1 87.0 91.0 76.4
CL+AR 70.4 44.4 79.8 81.5 86.4 70.9 85.0 89.4 74.6
Meta Llama-3.1 8B ZS 60.2 59.9 60.6 73.2 80.4 57.8 9.1 65.0 66.9 62.9 12.3
CoT 56.3 56.3 56.4 73.4 81.2 54.6 12.1 60.6 66.0 53.2 12.5
2S (Easy) 69.2 44.0 78.7 75.6 82.6 59.6 8.9 59.7 68.2 45.1 16.7
2S (Hard) 68.3 53.4 76.0 73.7 81.9 52.5 8.3 59.7 66.4 49.8 12.6
4S 71.6 50.9 80.0 76.4 83.3 60.1 6.9 59.9 66.7 49.7 11.7
AR 65.9 60.3 70.1 64.0 75.0 35.7 69.0 78.9 41.5
CL+AR 60.0 60.4 67.1 66.5 78.3 26.4 67.0 79.2 19.5
Llama-3.3 70B ZS 72.7 38.5 82.4 83.0 87.4 74.2 6.2 77.9 76.8 78.9 7.8
CoT 74.0 49.7 82.5 81.7 86.2 73.0 6.0 76.3 76.0 76.7 7.7
2S (Easy) 73.1 36.4 82.9 81.7 85.5 75.3 6.2 76.9 77.9 75.8 8.7
2S (Hard) 74.1 43.0 83.3 82.6 86.2 76.4 4.0 77.8 77.1 78.4 5.8
4S 73.2 40.4 82.7 81.9 85.5 75.8 4.7 78.7 78.4 79.0 6.7
AR 70.4 28.6 81.3 84.5 88.0 78.0 84.0 88.1 75.8
CL+AR 72.6 43.1 82.0 85.5 89.5 76.8 77.0 83.1 67.6
Llama-3.1 405B ZS 71.2 30.1 81.8 81.5 86.3 71.4 6.2 78.7 77.0 80.2 6.1
CoT 72.1 35.1 82.3 80.7 85.5 71.2 6.3 79.5 77.0 81.6 6.1
2S (Easy) 72.6 45.5 81.7 81.0 84.6 75.3 6.4 80.7 78.7 82.4 7.1
2S (Hard) 74.0 45.7 82.9 79.4 82.9 74.1 4.3 79.4 75.5 82.2 4.6
4S 75.2 51.0 83.4 79.6 82.9 74.8 5.6 81.0 78.4 83.1 5.8
AR 65.2 29.9 76.8 77.0 82.8 65.2 72.0 78.1 61.1
CL+AR 58.5 22.2 71.4 79.0 84.6 67.2 82.0 87.1 70.0
Disruption Monitor (8B) - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
Mistral Mixtral 8x7B ZS 70.5 42.4 80.2 58.4 56.1 60.5 11.9 62.5 37.4 73.2 13.5
CoT 67.0 21.5 79.1 57.9 54.6 60.8 13.2 64.4 40.4 74.6 13.9
2S (Easy) 69.7 30.0 80.7 64.9 65.6 64.2 10.3 67.4 54.3 74.7 11.9
2S (Hard) 71.5 50.8 79.9 68.0 69.5 66.4 6.3 66.8 52.2 74.5 10.1
4S 71.2 43.0 80.8 68.1 69.9 66.0 8.3 68.6 55.0 75.9 11.4
AR 65.2 43.6 75.1 58.5 66.9 44.3 60.0 69.7 41.2
CL+AR 65.9 44.4 79.8 62.5 70.1 49.7 62.0 68.9 51.3
Mixtral 8x22B ZS 70.7 28.5 81.6 81.8 85.9 74.7 8.9 74.3 64.8 79.8 14.3
CoT 70.8 33.4 81.3 80.9 85.4 72.5 9.0 73.7 64.0 79.3 14.7
2S (Easy) 73.8 55.8 81.4 83.5 87.4 76.0 6.7 78.9 75.8 81.3 9.9
2S (Hard) 73.2 61.7 79.4 81.1 84.9 74.5 4.2 75.5 69.9 79.4 6.6
4S 75.2 60.4 82.0 81.9 85.7 75.5 5.9 76.2 71.0 79.8 9.5
AR 68.1 31.7 79.2 81.0 86.8 66.1 79.0 84.2 68.7
CL+AR 63.7 47.3 72.3 75.5 84.0 47.3 83.0 88.3 69.1
DeepSeek DeepSeek-R1 ZS 73.8 55.9 81.3 81.1 86.4 69.1 6.7 74.5 76.7 71.8 8.8
CoT 74.9 57.5 82.2 80.4 86.0 67.6 6.6 76.6 78.1 74.9 7.6
2S (Easy) 75.8 57.8 83.0 82.3 87.0 72.5 6.4 72.7 75.6 69.0 10.1
2S (Hard) 76.4 65.4 82.1 82.0 86.9 71.4 4.4 72.5 76.1 67.7 6.9
4S 75.8 61.7 82.3 83.0 87.2 74.5 4.5 74.7 77.1 71.8 7.7
AR 71.1 60.6 77.2 80.0 86.2 63.6 85.0 90.2 68.1
CL+AR 75.6 66.7 80.7 80.0 86.5 61.5 87.0 91.4 73.5
Table 2: Experimental results of proprietary and open-source models on dialogue breakdown datasets. MSE represents verbalized confidence mean-squared error. Bold = best, underlined = second‐best.
6.1 Main Findings
State-of-the-Art Results. Our results confirm significant progress beyond earlier benchmarks. On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2). For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%). This tie represents the new top performance on DBDC5 English. On the DBDC5 Japanese track, the best outcome—89.0% accuracy—comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%. Thus, for Japanese, the Anthropic model retains a slight edge. Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
Closed-Source Frontier Models. Claude-3.5 (Haiku, Sonnet v2) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English. In particular, Sonnet v2 combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese. GPT-4o likewise competes closely, achieving up to 77.7% on BETOLD via 4S prompting and 83.5% on DBDC5 English using a more challenging 2S-Hard strategy. While marginally behind Claude-3.5 Sonnet v2 on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD and has the greatest accuracy swings falling near 41%-64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
Open-Source Models. Larger open-source models (Llama-3.3 70B, Llama-3.1 405B, Mixtral 8x22B, DeepSeek-R1) match or exceed closed-source baselines on DBDC5 English (80%-85%) and demonstrate competitive performance on DBDC5 Japanese. However, on BETOLD, performance variability is pronounced, ranging from 68.3% (Llama-3.1 8B, 2S-Hard) to 75.8% (DeepSeek-R1, 2S-Easy), reflecting difficulties generalizing from natural dialogues to abstract intent representations. Llama-3.1 405B, despite its size, does not consistently surpass the 70B model. Its best English-track accuracy hovers near 79%–81%, indicating marginal variances to prompting techniques.
Dialogue Disruption Monitor. Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance. On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models. For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes. Even without fine-tuning specifically for the BETOLD dataset, the model demonstrates improved generalization for dialogue breakdown detection, increasing accuracy from 60.2% in a ZS-baseline to 67.2%. On DBDC5 English, the model’s performance is comparable to its significantly larger variants while demonstrating better calibration scores. However, performance on the Japanese dataset remains limited due to insufficient Japanese training data. Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [45].
6.2 Impact of Prompting Strategies
Few-Shot Prompting. Few-shot prompting consistently outperforms ZS and CoT approaches across datasets. For instance, DeepSeek-R1 rises from 81.1% (ZS) to 83.0% (4S). The “Hard” exemplars comprised of borderline dialogues also yield stronger calibration (lower MSE). 2S-Hard achieves substantial calibration improvements, with Llama-3.3 70B reaching an MSE of 4.0 on DBDC5 English, the lowest among all models. Similarly, GPT-4o’s accuracy improves from 81.4% (ZS) to 83.5% (2S-Hard), affirming the effectiveness of providing challenging examples to refine model uncertainty estimations more effectively than simpler or even more examples.
Limitations of Chain-of-Thought. CoT yields mixed results, slightly improving borderline-case identification. For instance, GPT-4o on DBDC5 English F1(B) improves from 85.9% to 86.5%. However, this occasionally degrades performance on structured datasets like BETOLD, indicating an optimal reasoning complexity is dataset-dependent. This shows the importance of calibrating the number of reasoning steps: dialogues with short turns and domain-specific structure can be more efficiently parsed with concise prompts than with extensive “think step-by-step” sequences.
Challenges with AR and CL+AR on BETOLD. While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD. Results for BETOLD (highlighted in grey in Table 2) consistently showed lower performance. We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances. This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue. Through comprehensive error analysis, we observed that models frequently did not fully adhere to provided instructions. Common issues included either failing to generate any analogous examples—opting instead to solve the original dialogue directly—or producing overly brief summaries of analogous conversations that did not align well with the dialogue under evaluation. As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions. This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns). Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3. Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
6.3 Calibration and Confidence Analysis
Superior calibration (low MSE) indicates alignment between model confidence and human correctness. Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard). Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard). Meanwhile, our fine-tuned 8B monitor yields an MSE of 4.9, which is slightly higher but still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
6.4 Practical Costs and Sustainability
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[46]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 317 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 4,067 times higher [47]. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency—each dialogue inference completes in under 0.5 seconds. This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance. Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to slash costs by more than half compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations). Such selective invocation lowers operational costs and aligns with sustainability objectives. Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S. households [48]. By limiting resource-intensive model invocations to cases of genuine ambiguity or critical dialogue breakdowns, our approach effectively balances performance, cost-efficiency, and environmental sustainability.
7 Conclusion
Research and development in conversational AI robustness is rapidly advancing on multiple fronts. Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively. By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation. While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains. Short, curated exemplars typically offer a good trade-off between performance and token usage. The cost and energy savings are considerable, especially in frequent-turn applications. Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness. Our experiments reveal that carefully structured prompts—where the model justifies and then decides—yield lower calibration error. For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability. While closed-source models still lead in absolute accuracy, the performance gap is closing—especially once small and midsize open-source models are carefully tuned. Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
8 Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting stronger multilingual pretraining is needed. Moreover, real-world dialogues (involving code-switching, adversarial inputs, or varied domains) may demand specialized adaptations beyond our well-defined benchmarks, while prompt engineering remains model- and dataset-specific, increasing resource demands. The opaque nature of LLMs, even under chain-of-thought or analogical prompting, leaves internal reasoning partially hidden; consequently, escalating errors to more costly, higher-capacity models adds latency, energy use, and expense. Finally, simpler binary breakdown labels may miss borderline cases, pointing to a need for finer-grained schemes and more resource-aware architectures for robust deployment in diverse, high-stakes environments. Nonetheless, these constraints do not undermine the contributions of this work; we leave their resolution to future research.
Appendix A Cost-Savings Example on AWS Bedrock
To illustrate how real-time monitoring reduces expenses, we assume a 15-turn conversation where a user interacts primarily with Llama-3.3 70B but escalates to Llama-3.1 405B on 10% of turns if a breakdown is detected. Table 3 shows prices [^3] https://aws.amazon.com/bedrock/pricing/ per 1,000 tokens for Meta’s Llama-3.1 family.
Table 3: Selected AWS Bedrock prices for Llama 3.1.
Model Price / 1K Input Price / 1K Output
Llama-3.1 8B \$0.00022 \$0.00022
Llama-3.1 70B \$0.00072 \$0.00072
Llama-3.1 405B \$0.0024 \$0.0024
Assumptions
  • On average, each turn has 160 characters444www.twilio.com/docs/glossary/what-sms-character-limit (approximately 40 tokens) in the user prompt plus 40 tokens in the assistant’s reply, for 80 tokens total per turn.
  • At turn $i$, the LLM sees the entire conversation so far—so the token count accumulates.
  • Only 10% of turns require escalation to the 405B model.
Calculating Total Tokens Over 15 Turns
Since each turn $i$ includes all previous turns in the prompt, the approximate token usage over 15 turns:
$80\times\sum_{i=1}^{15}i=80\times\frac{15\times(15+1)}{2}=80\times 120=9600{}\text{tokens}$
Cost for Always Using 405B
  • Price: \$0.0048 per 1,000 tokens.
  • Total: $9.6\times\$0.0048\approx\$0.04608$
Cost for Always Using 70B
  • Price: \$0.00144 per 1,000 tokens.
  • Total: $9.6\times\$0.00144\approx\$0.01382$
Selective Escalation (70B + 10% 405B)
Assuming 10% of the total tokens trigger a breakdown and invoke the 405B model, while the remaining 90% stay on 70B:
  • 70B portion (8640 tokens): $8.64\,\text{k}\times\$0.00144=\$0.01244$
  • 405B portion (960 tokens): $0.96\,\text{k}\times\$0.0048=\$0.00460$
  • Combined: \$$\$0.01244+\$0.00460\approx\$0.01704$
Adding a Monitor for Real-time Checks
Every turn is first checked by a monitor:
  1. 1.
    Monitoring Cost:
    • 9,600 tokens at \$0.00044 per 1,000 tokens
    • $9.6\,\text{k}\times\$0.00044\approx\$0.00422$
  2. 2.
    Main Dialogue:
    • Still 90% on 70B and 10% on 405B \$$\approx\$0.01704$
  3. 3.
    Total: \$$\$0.00422+\$0.01704=\$0.02126$
Hence, even with monitoring overhead, escalating only 10% of turns to the 405B model cuts costs by more than half compared to always using Llama-3.1 405B.
Appendix B Additional Figures
Figure 3: BETOLD: CL+AR Prompt
Figure 4: Error Analysis: GPT-4 Analogical Reasoning Example on BETOLD